OpenStudy (anonymous):

Absolute min and max from f(t) = 2 cos t + sin 2t, [0, π/2]

zepdrix (zepdrix):

Mmm ok, find a derivative yet?

OpenStudy (campbell_st):

graphing it may be an easy solution.

OpenStudy (campbell_st):

use desmos https://www.desmos.com/calculator

OpenStudy (anonymous):

-2sint+cost ?

zepdrix (zepdrix):

Hmmmm not quite :O

OpenStudy (anonymous):

:(

zepdrix (zepdrix):

\[\Large\rm (\sin2t)'=(\cos2t)\color{royalblue}{(2t)'}\]\[\Large\rm (\sin2t)'=(\cos2t)\color{orangered}{(2)}\]Chain rule, yah? :o

OpenStudy (anonymous):

why are we using chain rule ?

zepdrix (zepdrix):

Wut? 0_o Because ... we ... always... use... chain rule :o

OpenStudy (anonymous):

oh sorry, I thought we always used it when there was a power ...Like (3x+5y)^2

zepdrix (zepdrix):

Chain rule is always necessary when the `inner function` is more than just "x". Woops, "t" in this case. Example:\[\Large\rm \frac{d}{dt}(t)^2=2(t)\color{royalblue}{\frac{d}{dt}(t)}=2(t)\color{orangered}{(1)}\]In this example, notice that by applying the chain rule, we gain nothing new? Yah it's not necessary here. But if the inner function was more than just t, Example:\[\Large\rm \frac{d}{dt}(3t)^2=2(3t)\color{royalblue}{\frac{d}{dt}(3t)}=2(3t)\color{orangered}{(3)}\]

zepdrix (zepdrix):

In our problem: Since our inner function was 2t instead of t, we needed to chain rule.

zepdrix (zepdrix):

Chain rule is by far the hardest of the quick shortcut rules to master :( Spend some time practicing it.

OpenStudy (anonymous):

oh alright, Soooo then sin2t = cost(2) would 2cost be -2sint ?

zepdrix (zepdrix):

The derivative of the first term looks good. D 2cos(t) = -2sin(t). We're still having some trouble with the other one though :(

zepdrix (zepdrix):

Ignore the chain rule for just a sec,\[\Large\rm \frac{d}{dt}\sin(2t)=\cos(2t)\]Yes? For some reason your 2t keeps turning into a t.

zepdrix (zepdrix):

If we properly apply our chain rule:\[\Large\rm \frac{d}{dt}\sin(2t)=\cos(2t)(2t)'\] \[\Large\rm \frac{d}{dt}\sin(2t)=2\cos(2t)\]Chain rule has no effect on our inner function. It shouldn't change it. We're making a copy of it, and multiplying by the derivative of that copy on the outside.

OpenStudy (anonymous):

Oops sorry. Then it would be = -2sin(t) + 2cos(2t) ?

zepdrix (zepdrix):

Mmm ok looks good,\[\Large\rm f'(t)=-2\sin (t)+2\cos(2t)\]

zepdrix (zepdrix):

Looking for critical points:\[\Large\rm 0=-2\sin (t)+2\cos(2t)\]

zepdrix (zepdrix):

It's going to be a little tricky from here. Do you remember your `Cosine Double Angle Identity`? It shows up in three different forms.

zepdrix (zepdrix):

This is the one that we want:\[\Large\rm \cos(2t)=2\cos^2(t)-1\]

OpenStudy (anonymous):

Oh yeah I remember that.

zepdrix (zepdrix):

Woops my bad, we have a sin(t) in our problem. I wrote the wrong one :( That's not the one we want.

zepdrix (zepdrix):

This is the one we want: \[\Large\rm \color{orangered}{\cos(2t)=1-2\sin^2(t)}\]

OpenStudy (anonymous):

aha thats what I wrote :)

zepdrix (zepdrix):

\[\Large\rm 0=-2\sin (t)+2\color{orangered}{\cos(2t)}\]hehe

zepdrix (zepdrix):

So that will give us a bunch of sin(t)'s which will make it easier to work with.

OpenStudy (anonymous):

ohh okay !

zepdrix (zepdrix):

\[\Large\rm 0=-2\sin (t)+2\color{orangered}{\left[1-2\sin^2(t)\right]}\]

OpenStudy (anonymous):

let x = sin ?

zepdrix (zepdrix):

Distribute some stuff... Divide a -2 out of everything... eventually we get to:\[\Large\rm 0=2sin^2(t)+\sin(t)-1\]A substitution? :) Mmmm yah that's a good idea!

zepdrix (zepdrix):

Leading to something like this, yah? :o \[\Large\rm 0=2x^2+x-1\]

zepdrix (zepdrix):

I skipped a few algebra steps in there >.< I think you can handle those hehe

zepdrix (zepdrix):

Solve for x using your quadratic formula, yah?

zepdrix (zepdrix):

brb i gotta get some chocolate milk all up in mah belly

zepdrix (zepdrix):

where you at lady? :3 figure it out?

OpenStudy (anonymous):

Sorry I fell asleep ! Ill look at it right now

OpenStudy (anonymous):

Got 1/2 and -1

OpenStudy (anonymous):

\[2x^2+2x-x-1=0,2x \left( x+1 \right)-1\left( x+1 \right)=0,\left( x+1 \right)\left( 2x-1 \right)=0\] x=-1,1/2 is correct

OpenStudy (anonymous):

\[\sin t=-1=-\sin \frac{ \pi }{ 2 }=\sin \left( \pi+\frac{ \pi }{ 2 } \right)=\sin \frac{ 3 \pi }{ 2 }\] \[t=\frac{ 3 \pi }{ 2 } does \not belong \to \left[ 0,\frac{ \pi }{ 2 } \right]\] hence rejected. \[\sin t=\frac{ 1 }{ 2 }=\sin \frac{ \pi }{ 6 },t=\frac{ \pi }{ 6 }\in \left[ 0,\frac{ \pi }{ 2 } \right]\] find f(t) at t=0,pi/6,pi/2 and see which is max. and which is min.